OpenStudy (anonymous):

You roll a standard number cube. Find P( number less than 3) Is it A? A) 1/3 B) 2/3 C) 3/2 D) 5/6

OpenStudy (rational):

numbers less than 3 : 1, 2

OpenStudy (rational):

so P(number less than 3) = 2/6 = 1/3 A is right !